OpenStudy (anonymous):

find the x-intercept of 4x-3y=-12

OpenStudy (therealmeeeee):

ok x-intercept is point where y=0: 4x-3(0)=12 4x=12 x=3 (3,0) y-intercept is point where x=0: 4(0)-3y=12 -3y=12 y=-4 (0,-4)

OpenStudy (therealmeeeee):

this helps

OpenStudy (anonymous):

yes, thank you. I missed a step.

OpenStudy (therealmeeeee):

medal

OpenStudy (therealmeeeee):

click Best Response

OpenStudy (anonymous):

Can I ask another question?

OpenStudy (therealmeeeee):

go ahead

OpenStudy (anonymous):

How would you divide 5m to the 3rd expo-9m to the 2nd expo+10m /5m to the 2nd expo?

OpenStudy (therealmeeeee):

wait

OpenStudy (therealmeeeee):

i don t know

OpenStudy (anonymous):

Ok, thanks!

OpenStudy (therealmeeeee):

click Best Response

OpenStudy (anonymous):

\[\frac{ 3 }{ x ^{2}? }+\frac{ 2 }{ x? }=1\] solve for x

OpenStudy (therealmeeeee):

do u understand

OpenStudy (anonymous):

no

OpenStudy (therealmeeeee):

oh thaats the best way i can put it im srry

OpenStudy (therealmeeeee):

and thats how i got it right

OpenStudy (anonymous):

I sent two problems. Which one are you talking about?

OpenStudy (therealmeeeee):

the first one

OpenStudy (anonymous):

Oh, yes I understand it. I just missed a step. Thanks for your help.